# Price predictability at ultra-high frequency: Entropy-based randomness test

### v2

@@ Abstract
We use the statistical properties of Shannon entropy estimator and Neyman-Pearson statistics to study the predictability of ultra-high frequency financial data. We develop a statistical test for the predictability of a sequence based on empirical frequencies. We study stylized facts that cause price predictability such as persistence of order signs, autocorrelation of returns, and volatility clustering. We show that the degree of randomness grows with the increase of aggregation level in transaction time. We also find that predictable days are usually characterized by high trading activity, i.e., days with unusually high trading volumes and the number of price changes. We find a group of stocks for which predictability is caused by a frequent change of price direction. We perform multiple testing for sub-intervals of days to identify whether there is predictability at a specific time period during the day.
Keywords test for predictability, limit order book, ultra-high frequency, entropy, empirical frequencies, Efficient market hypothesis
@@ 1 Introduction
One of the fundamental questions in finance is the predictability of asset prices. Currently, there is a set of tools designed to predict prices in the markets. For instance, traders rely on technical analysis of stocks to make a profit. The prediction power of technical analysis has been the object of many investigations, starting with the pioneering work [1]. We refer to [2] for a review of academic studies on this subject until 2007. For example, it has been found to generate excess profits in currency markets [3] including cryptocurrencies [4]. Profitable trading strategies can be constructed with the usage of forecasting methods such as neural networks [5]. Various trading methods can be developed in order to increase profits starting from prediction of future prices which are even only slightly more accurate than a martingale. For instance, an automatic trading method proposed in [6] outperforms a range of alternative approaches [^1] According to [7], abnormal returns based on price behavior in the market decline after academic publications about the profitable strategies associated with this behavior. It is indeed natural that the degree of market efficiency can change with time and is related to environmental factors characterizing “market ecology”, such as the number of competitors in the market, profit opportunities, and how the market participants adapt to changing environments. Under the adaptive market hypothesis [8], much of what behavioral finance cites as counterexamples to economic rationality is consistent with the evolutionary principles of financial interactions: competition, adaptation and natural selection; see [9, p. 7]..
@@
Randomness of financial data aggregated to daily frequency was investigated in the literature [10, 11, 12]. Studies analyzing the predictability of intraday prices were conducted using hourly [13, 14] and minute [15, 16] frequencies. Prices at millisecond and second frequencies were analyzed in [17]. This research progresses further toward a microscopic examination of financial time series. Ultra-high frequency data is defined as the full record of transactions and their associated characteristics. This research is dedicated to the predictability of ultra-high frequency data.
@@
Long memory of price returns at ultra-high frequency is a stylized fact that incorporates predictability into the time series. Lillo and Farmer [18] conducted a statistical test, concluding that both the signs of market orders and executed limit orders exhibit long-memory processes. They attributed this long memory of order signs to news arrivals and order splitting, which is offset by fluctuations in market liquidity. Therefore, the long memory does not contradict efficiency of markets when prices incorporate all available information about future values [19]. Moreover, the high speed of occurrence of new orders makes it difficult to predict the next price before it appears in such a short period of time. Without taking into account transaction costs, we can not ensure that predictability at ultra-high frequency indicates the presence of profitable trading strategies. However, we investigate the level of predictability as a function of the length of steps in transaction time. Then, we examine the periods in which the predictability of prices is presented. We explore what price characteristics distinguish days with predictability from others. For instance, we show that the autocorrelation of price returns is statistically significant during predictable days. Moreover, for most stocks, we observe a high probability of consecutive price directions across several transactions, aligning with the long-memory characteristics of price return signs.
@@
We devise a test for randomness of data starting from the estimation of Shannon entropy. Entropy is defined as an averaged measure of uncertainty about a symbol appearing in a sequence generated by a random source [20]. Maximum uncertainty arises when all symbols from a finite alphabet are independently generated with equal probabilities. A common method for entropy estimation is calculating empirical frequencies of blocks of symbols [21]. Previous instances where Shannon entropy served as a measure for assessing randomness in financial time series include [10, 12, 16]. The entropy of price returns of assets traded on the Moscow Stock Exchange was studied in [22]. The time-varying entropy of meme stocks, which experienced sudden price surges driven by social media communities, was investigated in [23]. In case of independent and identically distributed symbols, the difference between entropy estimation and its maximum follows $\chi^{2}$-distribution [24]. Using quantiles of the asymptotic $\chi^{2}$-distribution, we can determine significantly low values of entropy fast. We make a step forward and use a modification of entropy estimation that allows us to test uncertainty for a symbolic sequence in case when probabilities of symbols may be not equal. The test statistics is based on all overlapping blocks of symbols and takes into account their dependence. Calculating the statistic is straightforward, requiring only the estimation of frequencies for preselected block lengths.
@@
In a recent work [25], randomness of financial prices at daily frequency was tested using conditional probabilities. The authors tested if probabilities of price increase and decrease are the same conditionally on price history. We provide the test that considers that probabilities that a price moves up or down can be different but do not depend on previous history. We test if past price changes are helpful in predicting future price movement. To estimate Shannon entropy, we utilize discretization that distinguishes between positive and negative returns. A test for independence related to permutation entropy of price increments was introduced in [26]. Considering blocks of $k$ prices, frequencies of $k!$ events need to be calculated. In contrast, our statistic requires the calculation of only $2^{k}$ frequencies. Approximate entropy was used in [27, 28] to measure irregularity of financial time series. The authors fixed the length of blocks $k=3$. We use a larger length for the analysis based on a given sequence length. Entropy of a singular value decomposition was applied to test market efficiency in [29]. Monte Carlo simulations were employed by the authors to establish confidence intervals for low entropy values. In our test, we evaluate whether new outputs remain independent of the sequence history, leveraging statistics with known asymptotic distributions. Applying the test based on empirical frequencies of price directions, we get rid of Monte-Carlo simulations.
@@
This article introduces four key contributions. First, we propose the test for randomness of a symbolic sequence. It can be applied with varying numbers of distinct symbols, presuming the sequence length is sufficiently large. Additionally, the individual probabilities associated with each symbol’s appearance need not be equal. Second, we investigate the predictability of tick-by-tick data. We show that the degree of predictability decreases when we aggregate simulated and real data by a number of transactions. Third, we investigate the empirical properties of price returns and the difference in the properties of returns for not predictable and predictable time series. Some of the properties are stylized facts of price returns such as fat tails, autocorrelation of returns, and price jumps. We reveal that the probability of consecutive returns displaying the same sign (persistence) serves as a defining feature characterizing predictable sequences. For a group of stocks, this probability is significantly higher during predictable days. However, we also observe that the probability of repeating symbols is statistically low for predictable days in some stocks. For most stocks, we also find a high autocorrelation value inherent in predictable days. Other properties indicative of predictable days encompass heightened daily trading volumes and a substantial number of price changes.
Finally, we localize intervals that we call predictable and find the duration of periods of predictability. We employ the Šidák correction [30] to conduct multiple tests to assess price predictability within a single trading day. We detect intervals with different degrees of predictability inside a group of predictable trading days. Our analysis involves examining the duration of predictability periods in transaction time. We demonstrate that typically there exists a single predictable interval within a predictable day. It’s observed that two predictable intervals generally do not occur consecutively. In other words, following an interval with detected predictability, the subsequent interval doesn’t display a significant level of predictability. However, considering all transactions of the SNAP stock (Snap Inc.), we are able to detect several predictable time intervals going in a row.
In Section 2, we propose a statistical test for investigating the predictability of symbols of a sequence. Significance of values of entropy estimation and the Neyman-Pearson statistics are defined by $\chi^{2}$-distributions with suitable degrees of freedom. In Section 3, we present a chosen group of assets with a set of characteristics such as mean price, daily trading volumes, and number of transactions. Section 4.1 is dedicated to applying the predictability test to prices based on models involving long memory of order flows. In Section 4.2, we test the predictability of real data. We measure the predictability of daily time intervals while considering diverse block lengths, aggregation levels, and months. Section 4.3 explores the distinguishing factors between predictable and non-predictable sequences across various stocks and trading days. Section 4.4 investigates the predictability of pairs of price changes. For predictable days, we localize the presence of predictability inside predictable days in Section 4.5. Section 5 concludes the paper.
@@ 2 Statistical tests for randomness of symbolic sequence
The test is based on empirical frequencies of blocks of symbols introduced by Marton and Shields [21]. These empirical frequencies serve to estimate Shannon entropy, which is utilized as a measure of uncertainty. First, we divide a sequence by non-overlapping blocks with length $1<k<n$.
$\hat{X}={\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{n_{b}}}$
where $\hat{x}_{t}=\{x_{(t-1)k+1},x_{(t-1)k+2},\dots,x_{tk}\}$, $n_{b}=\lfloor\frac{n}{k}\rfloor$. $\hat{x}_{t}$ represents a number from 00 to $s^{k}-1$ in the numerical system with base $s$. We calculate $s^{k}$ empirical frequencies $\hat{f}_{j}$ of blocks from $\hat{X}$.
$\hat{f}_{j}=\sum_{t=1}^{n_{b}}I(\hat{x}_{t}=a_{j}),a_{j}\in A^{k}$
Here and further $I$ is an indicator function and $a_{j}\neq a_{l}$ when $j\neq l$. Then, entropy estimation is defined as
$\hat{H}=-\sum_{j=0}^{s^{k}-1}\frac{\hat{f}_{j}}{n_{b}}\ln{\frac{\hat{f}_{j}}{n_{b}}}$
where $\ln()$ is the natural logarithm with the convention $00\ln{0}=$ 0. According to [24] the entropy bias defined below (scaled distance from the maximum of entropy) follows $\chi^{2}$-distribution with $s^{k}-1$ degrees of freedom if the probabilities of all blocks of symbols, $\hat{x}_{t}$, are equal.
@@
When all probabilities of blocks of symbols are equal, the process is fully unpredictable and the entropy attains its maximum, $k\ln{s}$. We can test unpredictability of a sequence using the entropy bias and the distribution under the null hypothesis $\chi^{2}_{s^{k}-1}$. If the probabilities of appearing symbols (blocks with length 1) are not equal, then entropy estimation $\hat{H}$ has asymptotically normal distribution [24]. The mean and standard deviation of the normal distribution depend on the entropy value and probabilities of blocks, see [31, 32].
Now, we propose a statistics with overlapping blocks of symbols that we apply for testing predictability. For the test, we use the Neyman-Pearson criterion [33] that contains the likelihood of the sequence under the hypothesis of unpredictability. We use all blocks with length $10<k-1<n-$ 1.
$\bar{X}={\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{n-k+2}}$
@@
We construct statistics in Eq. 2 which has $\chi^{2}$-distribution under $\mathcal{H}_{0}$. The similar test statistics for Markov chains was considered by Billingsley [34].
$D=2\sum_{ij}f_{ij}\ln{\frac{(n-k+1)f_{ij}}{f_{\cdot j}f_{i\cdot}}}$ (2)
where $f_{ij}$ is the frequency of the event when a block with the last symbol $j\in A$ follows by the block numbered by $i$ in the sequence $\bar{X}$ and $f_{\cdot j}=\sum_{i}f_{ij}$, $f_{i\cdot}=\sum_{j}f_{ij}$; $10\leq i\leq M-$ 1, where $M=s^{k-1}$ is the amount of blocks of $k-1$ symbols. We note that $f_{ij}$ are also empirical frequencies of blocks of $k$ symbols. The convention is $00\ln{0}=$ 0 and $00\ln{(0/0)}=$ 0. The asymptotic distribution of the statistics $D$ is described in the Lemma provided below.
@@
Let’s consider a stationary process with symbols from a finite alphabet $A=\{0,1,\cdots,s-1\}$ and its realization $\{x_{t}\}_{t=1}^{n}$. The empirical frequencies of the k-blocks of the realization are as follows.
$f_{ij}=\sum_{t=1}^{n-k+1}I\left(\bar{x}_{t}=a_{i}\right)I\left(x_{t+k-1}=a_{j}\right),a_{i}\in A^{k-1},a_{j}\in A$
If the hypothesis of unpredictability, $\mathcal{H}_{0}$, holds true, then the Neyman-Pearson (NP) statistics (Eq. 2) converges in distribution to $\chi^{2}$ with $(s^{k-1}-1)(s-1)$ degrees of freedom, i.e.,
@@
We take $k=\left[0.5\log_{s}{n}\right]$ as the length of blocks that is proved to be admissible and suggested to use for estimating empirical frequencies in [35] [^2] See Sections 2.3.d and 3.2 or [21]. We provide the sketch of proof for the distribution of $D$ from Eq. 2 and subsequently verify it using Q-Q plots in Appendix Section A. We show that the proposed test statistics is valid even if the probabilities of appearing symbols are not equal. It is possible because of the denominator in Eq. 2 that mitigates differences in the probabilities of symbols inside the logarithm.
We calculate p-values associated with values of $B$ and $D$ and degrees of freedom of $\chi^{2}$-distribution. If a p-value is less than $0.01$, we conclude that the sequence does not exhibit significant predictability with the significance level $\alpha=0.01$. We reject $\mathcal{H}_{0}$ when p-value$<0.01$. We call intervals where $\mathcal{H}_{0}$ is rejected predictable.
The NP-statistics $D$ is similar to the Kullback–Leibler divergence [36] between empirical probabilities of blocks of symbols and the probabilities obtained under $\mathcal{H}_{0}$.
$\begin{split}&KL\left(\frac{f_{ij}}{(n-k+1)},\frac{f_{\cdot j}f_{i\cdot}}{(n-k+1)}\right)=\sum_{ij}\frac{f_{ij}}{n-k+1}\ln{\frac{(n-k+1)f_{ij}}{(n-k+1)f_{\cdot j}f_{i\cdot}}}\\ &=\frac{1}{(n-k+1)}\sum_{ij}f_{ij}\ln{\frac{(n-k+1)f_{ij}}{f_{\cdot j}f_{i\cdot}}}+\sum_{ij}\frac{f_{ij}}{n-k+1}\ln{(n-k+1)}\\ &=\frac{D}{2(n-k+1)}+\ln{(n-k+1)}\end{split}$

### v3

@@ Abstract
We use the statistical properties of Shannon entropy estimator and Kullback-Leibler divergence to study the predictability of ultra-high frequency financial data. We develop a statistical test for the predictability of a sequence based on empirical frequencies. We show that the degree of randomness grows with the increase of aggregation level in transaction time. We also find that predictable days are usually characterized by high trading activity, i.e., days with unusually high trading volumes and the number of price changes. We find a group of stocks for which predictability is caused by a frequent change of price direction. We study stylized facts that cause price predictability such as persistence of order signs, autocorrelation of returns, and volatility clustering. We perform multiple testing for sub-intervals of days to identify whether there is predictability at a specific time period during the day.
Keywords test for predictability, limit order book, ultra-high frequency, entropy, Kullback-Leibler divergence, Neyman-Pearson statistics, empirical frequencies
@@ 1 Introduction
One of the fundamental questions in finance is the predictability of asset prices. Currently, there is a set of tools designed to predict prices in the markets. For instance, traders rely on technical analysis of stocks to make a profit. The prediction power of technical analysis has been the object of many investigations [1, 2]. Various trading methods can be developed in order to increase profits starting from prediction of future prices which are even only slightly more accurate than a martingale [3, 4, 5, 6].
@@
Randomness of financial data aggregated to daily frequency was investigated in the literature [7, 8, 9]. Studies analyzing the predictability of intraday prices were conducted using hourly [10, 11] and minute [12, 13] frequencies. Prices at millisecond and second frequencies were analyzed in [14]. This research progresses further toward a microscopic examination of financial time series. Ultra-high frequency data is defined as the full record of transactions and their associated characteristics [15].
@@
This research is dedicated to the predictability of ultra-high frequency data. The high speed of occurrence of new orders makes it difficult to predict the next price before it appears in such a short period of time. Without taking into account transaction costs, we can not ensure that predictability at ultra-high frequency indicates the presence of profitable trading strategies. Moreover, a long memory of price returns at ultra-high frequency is a stylized fact that incorporates predictability into the time series but does not contradict the efficient market hypothesis [16, 17]. We study the relation between other stylized facts and predictability of financial time series at ultra-high frequency. This study complements the work [13] on stylized facts that add predictability to time series at a one-minute frequency. We investigate the level of predictability as a function of the length of steps in transaction time. By aggregating by the number of transactions, we increase the time between considered changes in price, dividing days into predictable and unpredictable days. We examine the periods in which the predictability of prices is presented. For instance, we observe for most stocks a high probability of consecutive price directions across several transactions, aligning with the long-memory characteristics of price return signs.
@@
We devise a test for randomness of data starting from the estimation of Shannon entropy. Entropy is defined as an averaged measure of uncertainty about a symbol appearing in a sequence generated by a random source [18]. Maximum uncertainty arises when all symbols from a finite alphabet are independently generated with equal probabilities. A common method for entropy estimation is calculating empirical frequencies of blocks of symbols [19]. Previous instances where Shannon entropy served as a measure for assessing randomness in financial time series include [7, 9, 13]. The entropy of price returns of assets traded on the Moscow Stock Exchange was studied in [20]. The time-varying entropy of meme stocks, which experienced sudden price surges driven by social media communities, was investigated in [21]. Such instabilities in a market and deviations from a fundamental price value was investigated in [22].
In case of independent and identically distributed symbols, the difference between entropy estimation and its maximum follows $\chi^{2}$-distribution [23]. We make a step forward and use the estimation of relative entropy that allows us to test uncertainty for a symbolic sequence in case when symbols can be not equiprobable. The test statistics is based on all overlapping blocks of symbols and takes into account their dependence. Calculating the statistic is straightforward, requiring only the estimation of frequencies for preselected block lengths. In a financial setting, a significantly high predictability implies that a future price change depends on the price history and thus is not a completely random. Empirical frequencies used in test statistics refer to patterns in sequences of price returns, some of which may be more likely to occur [20].
@@
In a recent work [24], randomness of financial prices at daily frequency was tested using conditional probabilities. The authors tested if probabilities of price increase and decrease are the same conditionally on price history. We provide the test for predictability that considers that probabilities that a price moves up or down can be different. We test if past price changes are helpful in predicting future price movement. We utilize discretization that distinguishes between positive and negative returns. A test for independence related to permutation entropy of price increments was introduced in [25]. Considering blocks of $k$ prices, frequencies of $k!$ events need to be calculated. In contrast, our statistic requires the calculation of only $2^{k}$ frequencies. Approximate entropy was used in [26, 27] to measure irregularity of financial time series. The authors fixed the length of blocks $k=3$. We use a larger length for the analysis based on a given sequence length. Entropy of a singular value decomposition was applied to test market efficiency in [28]. Monte Carlo simulations were employed by the authors to establish confidence intervals for low entropy values. In our test, we evaluate whether new outputs remain independent of the sequence history, leveraging statistics with known asymptotic distributions. Applying the test based on empirical frequencies of price directions, we get rid of Monte-Carlo simulations.
@@
This article introduces four key contributions. First, we propose the test for randomness of a symbolic sequence. It can be applied with varying numbers of distinct symbols, presuming the sequence length is sufficiently large. Additionally, the individual probabilities associated with each symbol’s appearance need not be equal. Second, we investigate the predictability of tick-by-tick data. We show that the degree of predictability decreases when we aggregate simulated and real data by a number of transactions. Third, we investigate the empirical properties of price returns and the difference in the properties of returns for not predictable and predictable time series. Some of the properties are stylized facts of price returns such as fat tails, autocorrelation of returns, and price jumps. Other properties indicative of predictable days encompass heightened daily trading volumes and a substantial number of price changes. We reveal that the probability of consecutive returns displaying the same sign (persistence) serves as a feature for predictable sequences. For a group of stocks, this probability is significantly higher during predictable days. On the other hand, we also observe that the probability of repeating symbols is statistically low for predictable days for another group of stocks. Finally, we localize intervals that we call predictable and find the duration of periods of predictability. We demonstrate that typically there exists a single predictable interval within a predictable day. It’s observed that two predictable intervals generally do not occur consecutively. In other words, following an interval with detected predictability, the subsequent interval doesn’t display a significant level of predictability. However, considering all transactions of the SNAP stock (Snap Inc.), we are able to detect several predictable time intervals going in a row. Analysis of the characteristics of predictability intervals and their duration for each stock provides insights into the price generation process.
In Section 2, we propose a statistical test for investigating the predictability of symbols of a sequence. In Section 3, we present a chosen group of diverse assets with a set of various characteristics. Section 4 is dedicated to applying the predictability test to simulated and real data. Section 5 concludes the paper.
@@ 2 Statistical tests for randomness of symbolic sequence
If the null hypothesis is rejected, the probability of guessing the next symbol of a sequence based on the last symbols, is higher than for a random guess. The test is based on empirical frequencies of blocks of symbols introduced by Marton and Shields [19]. These empirical frequencies serve to estimate Shannon entropy, which is utilized as a measure of uncertainty. The asymptotic distribution of the Shannon entropy estimation is known and presented in Lemma 1. The asymptotic distribution of the Neyman-Pearson statistics proposed to test predictability is given in Lemma 2.
First, we divide a sequence by non-overlapping blocks with length $k\in[2,n-1]$,
$\hat{X}={\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{n_{b}}},$
where $\hat{x}_{t}=\{x_{(t-1)k+1},x_{(t-1)k+2},\dots,x_{tk}\}$, $n_{b}=\lfloor\frac{n}{k}\rfloor$. $\hat{x}_{t}$ represents a number from 00 to $s^{k}-1$ in the numerical system with base $s$. We calculate $s^{k}$ empirical frequencies $\hat{f}_{j}$ of blocks from $\hat{X}$:
$\hat{f}_{j}=\sum_{t=1}^{n_{b}}I(\hat{x}_{t}=a_{j}),a_{j}\in A^{k}.$
Here and further $I$ is an indicator function and $a_{j}\neq a_{l}$ when $j\neq l$. The value of $k$ is user-defined. The upper bound of the length $k$ of symbolic patterns considered to find predictably is validated theoretically and empirically. We take $k=\left[0.5\log_{s}{n}\right]$ as the length of blocks that is proved to be admissible and suggested to use for estimating empirical frequencies in [29] [^1] See Sections 2.3.d and 3.2 or [19].
Then, entropy estimation is defined as
$\hat{H}=-\sum_{j=0}^{s^{k}-1}\frac{\hat{f}_{j}}{n_{b}}\ln{\frac{\hat{f}_{j}}{n_{b}}},$ (1)
where $\ln()$ is the natural logarithm with the convention $00\ln{0}=$ 0.
Lemma 1 (Asymptotic distribution of the Shannon entropy).
Let’s consider a stationary process with symbols from a finite alphabet $A=\{0,1,\cdots,s-1\}$ and its realizations $\{x_{t}\}_{t=1}^{n}$. Then, the entropy bias (scaled distance from the maximum of entropy) of a realization defined by Equations 1 and 2 follows $\chi^{2}$-distribution with $s^{k}-1$ degrees of freedom if the probabilities of all blocks of symbols, $\hat{x}_{t}$, are equal.
@@
See [23] for the proof. When all probabilities of blocks of symbols are equal, the process is fully unpredictable and the entropy attains its maximum, $k\ln{s}$. We can test unpredictability of a sequence using the entropy bias and the distribution under the null hypothesis $\chi^{2}_{s^{k}-1}$. If the probabilities of appearing symbols (blocks with length 1) are not equal, then entropy estimation $\hat{H}$ has asymptotically normal distribution [23]. The mean and standard deviation of the normal distribution depend on the entropy value and probabilities of blocks, see [30, 31].
Now, we propose a statistics with overlapping blocks of symbols that we apply for testing predictability. For the test, we use the Neyman-Pearson (NP) criterion [32] that contains the likelihood of the sequence under the hypothesis of unpredictability. We use all blocks with length $k-1$, $k\in[2,n-1]$,
$\bar{X}={\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{n-k+2}},$
@@
We construct statistics in Eq. 3 which has $\chi^{2}$-distribution under $\mathcal{H}_{0}$. The similar test statistics for Markov chains was considered by Billingsley [33]:
$D=2\sum_{ij}f_{ij}\ln{\frac{(n-k+1)f_{ij}}{f_{\cdot j}f_{i\cdot}}},$ (3)
where
$f_{ij}=\sum_{t=1}^{n-k+1}I\left(\bar{x}_{t}=a_{i}\right)I\left(x_{t+k-1}=a_{j}\right),a_{i}\in A^{k-1},a_{j}\in A$ (4)
is the frequency of the event when a block with the last symbol $j\in A$ follows by the block $a_{i}\in A^{k-1}$ in the sequence $\bar{X}$. We note that $f_{ij}$ are also empirical frequencies of blocks of $k$ symbols. $f_{\cdot j}=\sum_{i}f_{ij}$ and $f_{i\cdot}=\sum_{j}f_{ij}$, $10\leq i\leq M-$ 1, where $M=s^{k-1}$ is the amount of blocks of $k-1$ symbols. The convention is $00\ln{0}=$ 0 and $00\ln{(0/0)}=$ 0. The asymptotic distribution of the statistics $D$ is described in the Lemma provided below.
@@
Let’s consider a stationary process with symbols from a finite alphabet $A=\{0,1,\cdots,s-1\}$ and its realization $\{x_{t}\}_{t=1}^{n}$. If the hypothesis of unpredictability, $\mathcal{H}_{0}$, holds true, then the Neyman-Pearson (NP) statistics (Eq. 3, 4) converges in distribution to $\chi^{2}$ with $(s^{k-1}-1)(s-1)$ degrees of freedom, i.e.,
@@
We provide the sketch of proof for the distribution of $D$ from Equation 3 and subsequently verify it using Q-Q plots in Appendix Section A. We show that the proposed test statistics is valid even if the probabilities of appearing symbols are not equal. It is possible because of the denominator in Eq. 3 that mitigates differences in the probabilities of symbols inside the logarithm.
Significance of values of entropy estimation and the Neyman-Pearson statistics are defined by $\chi^{2}$-distributions with suitable degrees of freedom. We calculate p-values associated with values of $B$ and $D$ and degrees of freedom of $\chi^{2}$-distribution. If a p-value is less than $0.01$, we conclude that the sequence does not exhibit significant predictability with the significance level $\alpha=0.01$. We reject $\mathcal{H}_{0}$ when p-value$<0.01$. We call intervals where $\mathcal{H}_{0}$ is rejected predictable. While statistics $B$ is based on Shannon entropy estimation, statistics $D$ is scaled Kullback-Leibler (KL) divergence [34] between empirical probabilities $\frac{f_{ij}}{(n-k+1)}$ of blocks of symbols and the probabilities obtained under $\mathcal{H}_{0}$:
$\begin{split}&\text{KL}\left(\frac{f_{ij}}{(n-k+1)},\frac{f_{\cdot j}f_{i\cdot}}{(n-k+1)^{2}}\right)=\sum_{ij}\frac{f_{ij}}{n-k+1}\ln{\frac{(n-k+1)f_{ij}}{f_{\cdot j}f_{i\cdot}}}\\ &=\frac{1}{(n-k+1)}\sum_{ij}f_{ij}\ln{\frac{(n-k+1)f_{ij}}{f_{\cdot j}f_{i\cdot}}}=\frac{D}{2(n-k+1)}.\end{split}$ (5)
[…]